# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
Projection and economic hedging-demand recovery. Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (Stage 1+Stage 2) substantially reduces decision-time RMSE relative to the deployable Stage 1 policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, Stage 1 attains $3.54\text{\times}{10}^{-03}$ whereas Stage 1+Stage 2 achieves $1.56\text{\times}{10}^{-04}$. The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (Basic) and $1.42\text{\times}{10}^{-04}$ (Distill.) (Table 1), while the myopic RMSE is an order of magnitude smaller (Appendix Table 2). This pattern is consistent with the economic mechanism in this benchmark: once the (mostly) myopic component is captured, the dominant remaining challenge is to recover the intertemporal hedge induced by correlated return–factor shocks.
@@
Amortization, robustness, and the PPO baseline. Interactive distillation improves the deployable Stage 1 policy relative to the basic PG–DPO run, while the most accurate object remains the post-hoc projected policy (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This matches the intended division of labor in Section F: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
@@
As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables LABEL:tab:ou_rmse_s0_sweep_landscape1). Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 3).
@@
Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, and can be unstable in this terminal-only setting (Table LABEL:tab:ou_rmse_s0_sweep_landscape). This is consistent with PPO facing a difficult credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline. Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
@@ 6 Conclusion
We studied continuous-time portfolio choice in diffusion markets whose coefficients are estimated and therefore subject to statistical uncertainty (Section 2.1). We model this uncertainty by an exogenous law $q(d\theta)$ over a latent parameter $\theta$ that is drawn once at time 0 and remains fixed over the investment horizon, while the investor must deploy a single $\theta$-blind Markov feedback policy evaluated under an ex–ante CRRA objective (Remark 1, Section 2.1). This information structure shifts the relevant optimality notion from $\theta$-conditional (full-information) criticality to a $q$-aggregated Pontryagin first-order condition that is enforceable within the deployable $\theta$-blind policy class (Section 2.2, Theorem 1).
Methodologically, we extended Pontryagin–Guided Direct Policy Optimization (PG–DPO) to the latent-parameter setting by sampling $\theta$ only inside the simulator and computing exact discrete-time gradients via BPTT (Section 3.1), and we leveraged the BPTT–PMP correspondence to extract the costate objects needed for structured control updates (Theorem 2). Building on the $q$-aggregated stationarity, we proposed uncertainty-aware projected PG–DPO (P–PGDPO), which aggregates Monte Carlo Pontryagin quantities across $\theta\sim q$ and projects them onto the deployable first-order condition to obtain a single $\theta$-blind rule (Section 3.2). We established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3). In experiments, projection provides the dominant gains—improving stability and decision-time accuracy in high-dimensional drift-uncertainty scaling tests and recovering intertemporal hedging demand in factor-driven markets—while interactive distillation amortizes projection into a fast deployable network (Sections 4 and 5; Section F).
Several extensions are natural. A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to belief-aware decision rules (Remark 2, Appendix A). A second direction is to incorporate realistic frictions and constraints (transaction costs, leverage and short-sale limits) and develop certified or regularized projection steps when mixed-moment estimation becomes fragile (Section B; Section 4.2; Appendix E.5). Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
@@ Acknowledgments
Appendix A Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
@@
Purpose and scope. Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint. In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (Barberis, 2000; Pástor, 2000; Xia, 2001). This subsection records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter (a canonical partially observed diffusion setting; see, e.g., Bensoussan and van Schuppen (1985); Pham and Wei (2017)), and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$. We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization. In particular, our simulation-based Pontryagin-guided solvers developed later (Section 3) can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
@@
A linear–Gaussian hidden-premium model (OU state, observed returns). We use a stylized linear–Gaussian counterpart of the mean-reverting premium setting, but now assume the premium factor is not directly observed. Let $Y_{t}\in\mathbb{R}^{m}$ be a latent premium factor following an OU dynamics
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(\hat{y}_{0},P_{0}),$ (74)
@@
where $K\in\mathbb{R}^{m\times m}$ is stable, $\bar{y}\in\mathbb{R}^{m}$, and $\Xi\in\mathbb{R}^{m\times m}$. Risky assets satisfy
@@
We write $\mathbb{F}^{Z}=(\mathcal{F}_{t}^{Z})_{t\in[0,T]}$ for the filtration generated by $(Z_{s})_{s\leq t}$. For clarity, we present the independent-noise case $W\perp W^{Y}$; the correlated-noise extension remains linear–Gaussian but leads to more cumbersome gain formulas.
@@
$\dot{P}_{t}$ $=KP_{t}+P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (79)
Thus, even though the posterior $q_{t}$ is a distribution-valued object, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (79); this is the prototypical setting in which belief-state control reduces to finite-dimensional sufficient statistics (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization. To mirror the decision-time perspective of the OU benchmark, we consider the remaining-horizon time-averaged premium
$\bar{\theta}_{t,T}:=\frac{1}{T-t}\int_{t}^{T}BY_{s}\,ds\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (80)
For the OU dynamics (74), one has the decomposition
$\int_{t}^{T}Y_{s}\,ds=\tau\bar{y}+K^{-1}\big(I-e^{-K\tau}\big)(Y_{t}-\bar{y})+\int_{t}^{T}K^{-1}\big(I-e^{-K(T-u)}\big)\Xi\,dW_{u}^{Y}.$ (81)
Conditioning on $\mathcal{F}_{t}^{Z}$, the random variable $Y_{t}$ is distributed as $\mathcal{N}(\hat{Y}_{t},P_{t})$ by (77), while the future increments $(W_{u}^{Y}-W_{t}^{Y})_{u\geq t}$ are independent of $\mathcal{F}_{t}^{Z}$ in the independent-noise case. Hence $\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}$ is Gaussian:
$\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}\!\big(m_{t,T},\,P_{t,T}\big),\hskip 18.49988ptm_{t,T}:=\frac{B\,m_{I}(t,T)}{\tau},\hskip 18.49988ptP_{t,T}:=\frac{1}{\tau^{2}}\,B\,C_{I}(t,T)\,B^{\top},$ (82)
@@
Equation (82) provides a concrete example of an online-updated Gaussian uncertainty description $q_{t,T}:=\mathcal{L}(\bar{\theta}_{t,T}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(m_{t,T},P_{t,T}).$
@@
A plug-in decision-time benchmark (receding-horizon fixed-$q_{t,T}$). Given $(m_{t,T},P_{t,T})$ from (82), a simple decision-time rule is obtained by treating $q_{t,T}$ as fixed over the remaining horizon and applying the Gaussian constant-allocation benchmark of Section 2.3.1 with horizon $\tau$:
$\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,T}\Big)^{-1}m_{t,T}.$ (85)
@@
One may interpret (85) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the general “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (Barberis, 2000; Pástor, 2000; Xia, 2001).
@@
Remarks (relation to belief-aware control). The plug-in rule (85) is intentionally decision-time: it conditions on the current uncertainty description and does not attempt to optimize over how the posterior will evolve. In the present paper, we therefore treat the uncertainty law as fixed at a decision time (either as a fixed $q$ over a horizon, or as an externally updated sequence of inputs $q_{t}$ that is not controlled by the agent), mirroring the decision-time perspective common in Bayesian/learning portfolio-choice studies (Barberis, 2000; Pástor, 2000; Xia, 2001). Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (Haussmann, 1987; Li and Tang, 1995; Baghery and Øksendal, 2007). Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in $(x,y,\hat{Y},P)$-space (or its sufficient-statistic analogues) is an important direction that we defer to future work.

### v3

@@ 5.2 Results: hedging-demand recovery, amortization, and robustness to decision-time uncertainty
Projection and economic hedging-demand recovery. Across all $d$ and $s_{0}$, the post-hoc Pontryagin projection (PG–DPO) substantially reduces decision-time RMSE relative to the deployable Stage 1 warm-up (DPO) policy (Table 2). For example, under aligned $P_{0}$ with $s_{0}=10^{-3}$ and $d=100$, DPO attains $3.54\text{\times}{10}^{-03}$ whereas PG–DPO achieves $1.56\text{\times}{10}^{-04}$. The component-wise diagnostics indicate that the remaining discrepancy is largely driven by the hedging channel: in the same setting, the hedging RMSE is $1.55\text{\times}{10}^{-04}$ (basic) and $1.42\text{\times}{10}^{-04}$ (distill.) (Table 3), while the myopic RMSE is an order of magnitude smaller (Appendix Table 4). This pattern is consistent with the economic mechanism in this benchmark: once the (mostly) myopic component is captured, the dominant remaining challenge is to recover the intertemporal hedge induced by correlated return–factor shocks.
@@
Amortization, robustness, and the PPO baseline. Interactive distillation improves the deployable Stage 1 policy relative to basic Stage 1 (DPO), while the most accurate object remains the post-hoc projected policy (PG–DPO) (Table 2). This matches the intended division of labor in Section G: Stage 2 provides a structured stationarity-correction signal through the aggregated Pontryagin projection, and distillation amortizes that correction into a single forward pass, up to policy-class approximation limits.
@@
As the decision-time uncertainty scale $s_{0}$ increases, both the full RMSE and the hedging-component RMSE increase, with the most visible degradation at $s_{0}=10^{-1}$, especially at larger dimensions (Tables 23). Misalignment has a limited effect for small and moderate uncertainty scales, but can induce noticeable deterioration in the hardest settings, where the direction-of-hedge diagnostic can also weaken (Appendix Table 5).
$s_{0}$ Method $d=5$ $10$ $50$ $100$ Aligned $P_{0}$ $10^{-3}$ PG–DPO (basic) $4.59\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $10^{-3}$ PG–DPO (distill.) $4.39\text{\times}{10}^{-05}$ $5.25\text{\times}{10}^{-05}$ $1.37\text{\times}{10}^{-04}$ $1.42\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (basic) $98\text{\times}{10}^{-05} 4.4$$47\text{\times}{10}^{-05}$ $1.44\text{\times}{10}^{-04}$ $1.74\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (distill.) $4.43\text{\times}{10}^{-05}$ $4.95\text{\times}{10}^{-05}$ $1.47\text{\times}{10}^{-04}$ $1.55\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (basic) $5.27\text{\times}{10}^{-05}$ $3.99\text{\times}{10}^{-05}$ $2.60\text{\times}{10}^{-04}$ $2.95\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (distill.) $4.72\text{\times}{10}^{-05}$ $5.36\text{\times}{10}^{-05}$ $2.58\text{\times}{10}^{-04}$ $2.86\text{\times}{10}^{-04}$ Misaligned $P_{0}$ $10^{-3}$ PG–DPO (basic) $5.00\text{\times}{10}^{-05}$ $4.87\text{\times}{10}^{-05}$ $1.34\text{\times}{10}^{-04}$ $1.57\text{\times}{10}^{-04}$ $10^{-3}$ PG–DPO (distill.) $4.30\text{\times}{10}^{-05}$ $5.33\text{\times}{10}^{-05}$ $1.40\text{\times}{10}^{-04}$ $1.43\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (basic) $4.93\text{\times}{10}^{-05}$ $5.65\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.75\text{\times}{10}^{-04}$ $10^{-2}$ PG–DPO (distill.) $4.56\text{\times}{10}^{-05}$ $6.09\text{\times}{10}^{-05}$ $1.53\text{\times}{10}^{-04}$ $1.56\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (basic) $5.45\text{\times}{10}^{-05}$ $1.55\text{\times}{10}^{-04}$ $3.19\text{\times}{10}^{-04}$ $3.28\text{\times}{10}^{-04}$ $10^{-1}$ PG–DPO (distill.) $5.20\text{\times}{10}^{-05}$ $1.57\text{\times}{10}^{-04}$ $3.13\text{\times}{10}^{-04}$ $3.20\text{\times}{10}^{-04}$
Table 3: Decision-time RMSE at $t=0$ for the hedging component of the OU decision-time reference, evaluated on the projected (Stage 2 / PG–DPO) policies (tail median over the last six evaluation checkpoints). We use the decomposition $\pi^{\mathrm{ref}}(\tau)=\pi^{\mathrm{myo}}(\tau)+\pi^{\mathrm{hedge}}(\tau)$ where $\pi^{\mathrm{myo}}$ is the $\rho=0$ reference and $\pi^{\mathrm{hedge}}$ is the residual $\rho$-dependent term (cf. (37)). Component-wise diagnostics are reported for projected rules since Stage 1 (DPO) and PPO do not explicitly output a compatible myopic/hedging split under our protocol.
@@
Finally, the PPO baseline remains far from the analytic OU reference under the same $Y$-blind deployability restriction, and can be unstable in this terminal-only setting (Table 2). This is consistent with PPO facing a difficult credit-assignment problem under latent-factor heterogeneity, in contrast to the pathwise-sensitivity and affine-in-control correction exploited by our two-stage pipeline. Since PPO does not provide a compatible myopic/hedging decomposition under our evaluation protocol, we include it only in the full-RMSE table.
@@ 6 Conclusion
We studied continuous-time CRRA portfolio choice in diffusion markets whose coefficients are estimated and therefore statistically uncertain (Section 2.1). Our modeling choice is to treat estimation risk as an exogenous decision-time input, represented by a law $q(d\theta)$ over a latent parameter $\theta$: Nature draws $\theta$ once at time 0 and keeps it fixed over the investment horizon, while the investor does not observe $\theta$ and must deploy a single $\theta$-blind Markov feedback policy evaluated under an ex–ante objective (Remark 1). Under this information structure, the appropriate optimality notion is not the infeasible $\theta$-conditional (full-information) criticality, but rather a $q$-aggregated Pontryagin first-order condition that is enforceable within the deployable $\theta$-blind policy class (Section 2.2, Theorem 1).
Methodologically, we proposed a simulation-only two-stage solver, PG–DPO, that matches this deployability constraint. In Stage 1, DPO optimizes the ex–ante objective by sampling $\theta$ only inside the simulator and performing stochastic gradient ascent via exact discrete-time pathwise gradients computed by BPTT through an Euler discretization (Section 3.1). We then leverage a conditional BPTT–PMP correspondence to extract the costate objects needed for structured control updates, including the second-adjoint blocks entering the portfolio Hamiltonian gradient (Theorem 2). In Stage 2, we apply a $q$-aggregated Pontryagin projection: we aggregate Monte Carlo Pontryagin quantities across $\theta\sim q$ and project onto the deployable $q$-aggregated stationarity condition to obtain a single $\theta$-blind rule (Section 3.2). On the theory side, we established a residual-based ex–ante policy-gap bound under local stability of the aggregated projection map, with discretization and Monte Carlo errors made explicit (Theorem 3).
Empirically, across benchmarks the projection step provides the dominant gains, improving stability and decision-time accuracy in high-dimensional drift-uncertainty scaling tests and recovering intertemporal hedging demand in factor-driven markets, while interactive distillation amortizes the projected rule into a fast deployable network (Sections 4 and 5; Section G).
Several extensions are natural. A first direction is to allow time-varying uncertainty descriptions $q_{t}$ (e.g., produced by an external filter) and connect the present fixed-$q$ projection to decision-time plug-in replanning and belief-aware decision rules (Remark 3, Appendix B). A second direction is to incorporate realistic frictions and constraints—such as transaction costs, leverage, and short-sale limits—and develop certified or regularized projection steps for regimes where mixed-moment estimation becomes fragile (Section C; Section 4.2; Appendix F.5). Finally, applying the framework to large cross-sectional datasets with modern estimation pipelines would further clarify the practical benefits of inference-agnostic, simulation-only optimization under parameter uncertainty (Section 1).
@@ Acknowledgments
Appendix A Derivations for Gaussian decision-time references
This appendix derives the closed-form decision-time constant-portfolio references stated in Section 2.3. Throughout, we fix a decision time $t$ and hold the input law $q_{t}$ fixed over the remaining horizon $[t,T]$ (plug-in replanning).
A.1 Static latent drift: tilted FOC and Gaussian shrinkage
Consider the static latent-drift model (24)–(26) and restrict to constant portfolio fractions $\pi_{s}\equiv\pi$ on $[t,T]$. Let $\tau:=T-t$ and CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$.
Step 1: conditional evaluation under fixed $\theta$.
For constant $\pi$, conditional on $\theta$ the wealth satisfies
$\log\frac{X_{T}^{\pi}}{x}=\Big(r+\pi^{\top}\theta-\tfrac{1}{2}\pi^{\top}\Sigma\pi\Big)\tau+\pi^{\top}\Sigma^{1/2}(W_{T}-W_{t}),$
hence $X_{T}^{\pi}$ is lognormal and
$\mathbb{E}\big[U(X_{T}^{\pi})\mid\theta,\mathcal{F}_{t}\big]=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)r\tau+(1-\gamma)\tau\,\pi^{\top}\theta-\tfrac{1}{2}\gamma(1-\gamma)\tau\,\pi^{\top}\Sigma\pi\Big\}.$ (68)
Step 2: ex-ante mixing over $q_{t}$.
Mixing over $\theta\sim q_{t}$ yields
where $M_{q_{t}}(u):=\mathbb{E}_{\theta\sim q_{t}}[\exp(u^{\top}\theta)]$ is the mgf of $q_{t}$.
Step 3: tilted first-order condition.
Maximizing $J_{t}(\pi)$ is equivalent to maximizing $\log J_{t}(\pi)$. Differentiating (69) and using interior optimality gives the tilted FOC
$\gamma\,\Sigma\,\pi=\nabla_{u}\log M_{q_{t}}(u)\Big|_{u=(1-\gamma)\tau\,\pi}.$ (70)
Step 4: Gaussian input law.
If $q_{t}=\mathcal{N}(m_{t},P_{t})$, then $M_{q_{t}}(u)=\exp\{u^{\top}m_{t}+\tfrac{1}{2}u^{\top}P_{t}u\}$ and $\nabla_{u}\log M_{q_{t}}(u)=m_{t}+P_{t}u$. Substituting into (70) yields the linear system
$\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)\,\pi=m_{t},$ (71)
and hence
$\pi=\big(\gamma\Sigma-(1-\gamma)\tau\,P_{t}\big)^{-1}m_{t}=\big(\gamma\Sigma+(\gamma-1)\tau\,P_{t}\big)^{-1}m_{t},\hskip 18.49988pt(\gamma>1),$ (72)
which matches (30) in the main text.
A.2 OU premium factor: induced effective law and constant-portfolio reference
We derive the OU-based reference stated in (37). Fix $t$ and let $\tau:=T-t$. Under (31)–(33), we take as decision-time input $Y_{t}\sim\mathcal{N}(m_{t},P_{t})$ and hold this input fixed over $[t,T]$.
A.2.1 Integrated factor: mean and covariance
Define $I_{t,T}:=\int_{t}^{T}Y_{s}\,ds$. Assume $K$ is positive stable (equivalently, K-K is Hurwitz), so that $e^{-Ku}$ decays as $u\to\infty$. For $u\geq 0$ define
$\mathcal{A}(u):=K^{-1}\big(I_{m}-e^{-Ku}\big),$
where $I_{m}$ is the $m\times m$ identity matrix. The OU solution is
$Y_{t+u}=\bar{y}+e^{-Ku}(Y_{t}-\bar{y})+\int_{0}^{u}e^{-K(u-s)}\Xi\,dW_{t+s}^{Y}.$
Integrating over $u\in[0,\tau]$ gives
$I_{t,T}=\tau\bar{y}+\mathcal{A}(\tau)(Y_{t}-\bar{y})+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\,dW_{t+u}^{Y}.$
Taking expectation under the input law yields
$m_{I}(t,\tau):=\mathbb{E}[I_{t,T}]=\tau\bar{y}+\mathcal{A}(\tau)\,(m_{t}-\bar{y}),$ (73)
and the covariance decomposition gives
$C_{I}(t,\tau):=\mathrm{Cov}(I_{t,T})=\mathcal{A}(\tau)\,P_{t}\,\mathcal{A}(\tau)^{\top}+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\Xi^{\top}\,\mathcal{A}(u)^{\top}\,du.$ (74)
A.2.2 Effective premium law
Define the horizon-averaged effective premium
$\bar{\theta}_{t,\tau}:=\frac{1}{\tau}B\,I_{t,T}.$
Then $\bar{\theta}_{t,\tau}$ is Gaussian with
$m_{\bar{\theta}}(t,\tau)=\frac{1}{\tau}B\,m_{I}(t,\tau),\hskip 18.49988ptP_{\bar{\theta}}(t,\tau)=\frac{1}{\tau^{2}}B\,C_{I}(t,\tau)\,B^{\top},$ (75)
matching (36).
A.2.3 Cross term induced by return–factor correlation
Let $\Delta W:=W_{T}-W_{t}$. Using $d\langle W,W^{Y}\rangle_{s}=\rho\,ds$, we obtain
$C_{IW}(\tau):=\mathrm{Cov}\!\big(I_{t,T},\,\Delta W\big)=\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\,\rho^{\top}\,du\;\in\;\mathbb{R}^{m\times d},$ (76)
and the induced symmetric cross term
A.2.4 Constant-portfolio CRRA objective and FOC
Restrict to constant $\pi$ over $[t,T]$. Write the excess-return integral as
$\int_{t}^{T}\pi^{\top}dR_{s}=\pi^{\top}BI_{t,T}+\pi^{\top}\Sigma^{1/2}\Delta W.$
As in the static case, CRRA evaluation reduces to the log-mgf of a jointly Gaussian vector. Collecting the terms involving $\pi$ and applying the interior first-order condition yields the linear system
$\Big(\gamma\tau\Sigma+(\gamma-1)\big(B\,C_{I}(t,\tau)\,B^{\top}+M_{\mathrm{cross}}(\tau)\big)\Big)\,\pi=B\,m_{I}(t,\tau),\hskip 18.49988pt(\gamma>1),$ (78)
which matches (37) in the main text.
Independence case.
If $\rho=0$, then $C_{IW}(\tau)=0$ and $M_{\mathrm{cross}}(\tau)=0$. In this case, (78) reduces to the static Gaussian shrinkage formula applied to the effective law (75).
@@ A.2.4 Constant-portfolio CRRA objective and FOC
Appendix B Online uncertainty updates: Kalman–Bucy filtering and a plug-in decision-time benchmark
@@
Purpose and scope. Sections 2.3.1 and 2.3.2 focus on decision-time benchmarks in which an uncertainty description $q$ is treated as given and the investor optimizes under the corresponding $\theta$-blind deployability constraint. In practice, however, new data arrive and the uncertainty description is updated over time by an external estimation/filtering engine, a viewpoint that aligns with learning/estimation-risk portfolio choice and Bayesian decision-time formulations (Barberis, 2000; Pástor, 2000; Xia, 2001). This appendix records a simple linear–Gaussian example in which such an updated description $q_{t}$ arises endogenously via a Kalman–Bucy filter, and then formalizes a plug-in workflow: at each decision time, treat the current uncertainty description $q_{t}$ as given and compute a decision-time optimal control under that $q_{t}$.
We emphasize that solving the fully optimal partial-observation (belief-state) control problem is not the goal of this paper; rather, we view the resulting $q_{t}$ as an external input to decision-time optimization. In particular, the simulation-based Pontryagin-guided solvers developed in Section 3 can be used as inner-loop engines that are refreshed whenever a new uncertainty description $q_{t}$ becomes available.
@@
A linear–Gaussian hidden-premium model (OU state, observed returns). Let $Y_{t}\in\mathbb{R}^{m}$ be a latent premium factor following an OU dynamics
$dY_{t}=K(\bar{y}-Y_{t})\,dt+\Xi\,dW_{t}^{Y},\hskip 18.49988ptY_{0}\sim\mathcal{N}(m_{0},P_{0}),$ (79)
@@
where $K$ is positive stable (equivalently K-K is Hurwitz), $\bar{y}\in\mathbb{R}^{m}$, and $\Xi\in\mathbb{R}^{m\times m}$. Risky assets satisfy
@@
We write $\mathbb{F}^{Z}=(\mathcal{F}_{t}^{Z})_{t\in[0,T]}$ for the filtration generated by $(Z_{s})_{s\leq t}$.
For clarity, we present the independent-noise case $W\perp W^{Y}$; the correlated-noise extension remains linear–Gaussian but leads to more cumbersome gain formulas.
@@
$\dot{P}_{t}$ $=-KP_{t}-P_{t}K^{\top}+\Xi\Xi^{\top}-P_{t}B^{\top}\Sigma^{-1}BP_{t},\hskip 18.49988ptP_{0}\ \text{given}.$ (84)
Thus, even though $q_{t}$ is distribution-valued, in this affine/Gaussian regime it is fully characterized by the finite-dimensional sufficient statistics $(\hat{Y}_{t},P_{t})$, with $P_{t}$ evolving deterministically via (84).
From a posterior on $Y_{t}$ to a Gaussian uncertainty description for decision-time optimization. Define the remaining-horizon time-averaged premium
$\bar{\theta}_{t,\tau}:=\frac{1}{\tau}\int_{t}^{T}BY_{s}\,ds\in\mathbb{R}^{d},\hskip 18.49988pt\tau:=T-t.$ (85)
Let $\mathcal{A}(u):=K^{-1}(I_{m}-e^{-Ku})$, where $I_{m}$ is the $m\times m$ identity matrix. The integrated factor admits the representation
$\int_{t}^{T}Y_{s}\,ds=\tau\bar{y}+\mathcal{A}(\tau)\,(Y_{t}-\bar{y})+\int_{0}^{\tau}\mathcal{A}(u)\,\Xi\,dW_{t+u}^{Y}.$ (86)
Conditioning on $\mathcal{F}_{t}^{Z}$, we have $Y_{t}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}(\hat{Y}_{t},P_{t})$ by (82), and the future increments $(W_{t+u}^{Y}-W_{t}^{Y})_{u\geq 0}$ are independent of $\mathcal{F}_{t}^{Z}$ in the independent-noise case. Hence $\bar{\theta}_{t,\tau}\mid\mathcal{F}_{t}^{Z}$ is Gaussian:
$\bar{\theta}_{t,\tau}\mid\mathcal{F}_{t}^{Z}\sim\mathcal{N}\!\big(m_{t,\tau},\,P_{t,\tau}\big),\hskip 18.49988ptm_{t,\tau}:=\frac{B\,m_{I}(t,\tau)}{\tau},\hskip 18.49988ptP_{t,\tau}:=\frac{1}{\tau^{2}}\,B\,C_{I}(t,\tau)\,B^{\top},$ (87)
@@
Equation (87) thus provides an explicit example of an online-updated Gaussian uncertainty description $q_{t,\tau}:=\mathcal{L}(\bar{\theta}_{t,\tau}\mid\mathcal{F}_{t}^{Z})=\mathcal{N}(m_{t,\tau},P_{t,\tau}).$
@@
A plug-in decision-time benchmark (receding-horizon fixed-$q_{t,\tau}$). Given $(m_{t,\tau},P_{t,\tau})$ from (87), a simple decision-time rule is obtained by treating $q_{t,\tau}$ as fixed over the remaining horizon and applying the Gaussian constant-allocation benchmark of Section 2.3.1 with horizon $\tau$:
$\pi_{t}^{\mathrm{plug}}:=\Big(\gamma\,\Sigma+(\gamma-1)\tau\,P_{t,\tau}\Big)^{-1}m_{t,\tau}.$ (90)
@@
One may interpret (90) as a receding-horizon decision-time policy driven by an externally updated uncertainty description, consistent with the “update beliefs, then optimize” workflow used in Bayesian/learning-based portfolio choice (Barberis, 2000; Pástor, 2000; Xia, 2001).
@@
Remarks (relation to belief-aware control). The plug-in rule (90) is intentionally decision-time: it conditions on the current uncertainty description and does not optimize over how the posterior will evolve. Even in linear–Gaussian regimes where the belief state is finite-dimensional, the fully optimal partial-observation portfolio problem would treat the belief state (here, $(\hat{Y}_{t},P_{t})$) as part of the controlled state and optimize the policy in that enlarged state space (Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Related necessary conditions under partial information can also be expressed via partial-observation maximum principles (Haussmann, 1987; Li and Tang, 1995; Baghery and Øksendal, 2007). Developing a belief-aware Pontryagin-guided policy optimizer that operates directly in belief-state space is an important direction that we defer to future work.